# Examining Two Hop Reasoning Through Information Content Scaling

### v1

@@ Abstract
Prior work has found that transformers have an inconsistent ability to learn to answer latent two-hop questions – questions of the form “Who is Bob’s mother’s boss?” We study why this is the case by examining how transformers’ capacity to learn datasets of two-hop questions and answers (two-hop QA) scales with their size, motivated by prior work on transformer knowledge capacity for simple factual memorization. We find that capacity scaling and generalization both support the hypothesis that latent two-hop QA requires transformers to learn each fact twice, while two-hop QA with chain of thought does not. We also show that with appropriate dataset parameters, it is possible to “trap” very small models in a regime where they memorize answers to two-hop questions independently, even though they would perform better if they could learn to answer them with function composition. Our findings show that measurement of capacity scaling can complement existing interpretability methods, though there are challenges in using it for this purpose.
@@ 1.1 Motivation
Many existing transformer interpretability methods, like probing (Belinkov, 2022) and sparse autoencoders (Huben et al., 2023) mostly involve extracting interpretable signals from the activations. These allow us to identify features that are relevant to the computation for particular data points, but they also necessarily ignore other parts of the computation. Another ubiquitous method for understanding models is to test how they generalize to situations absent from the training data – this is extremely useful, but so standard it would not typically be considered an “interpretability technique”. Both methods come with limitations: hidden state probing may fail to capture important features of the computation or fail to yield useful hypotheses about how the model is computing answers. It is usually infeasible to test a model’s generalization performance on all tasks that might be of interest.
We investigate a new, complementary method for interpreting transformers: information content. Prior work by Allen-Zhu & Li (2024) found that sufficiently trained transformers can memorize facts up to a limit of around 2 bits per parameter, independent of architecture or dataset size. Let’s consider carefully what this figure might represent. It cannot represent the maximum rate of compression of every trained transformer. Transformers are feedforward networks, so if a dataset benefits from recurrence, the best a transformer can do is learn several layers that all apply the same function (perhaps up to some symmetry). A good compression algorithm could recognise the repetition of layers, and encode a single layer along with the instruction to repeat it a sufficient number of times, but this affords no opportunity to the transformer to pack the computation into fewer parameters.
@@
We hypothesize that the figure of 2 bits per parameter represents how a transformer can be compressed with respect to a model of computation that matches the transformer’s architecture. This compression is very efficient for factual recall tasks, because they do not benefit from recurrence, but will be less efficient in tasks that do benefit from recurrence. We will call this imprecise notion the “native information content” of a transformer, and we will study how this figure scales with model size for the task of two-hop question answering.
The native information content of transformers is interesting because it relates to how well transformers generalize at different tasks. Information compression is closely linked to learning in general – given a sequence of data $(X_{1},X_{2},...)$ sampled from a distribution $P(X)$, the we can can model $P$, the more efficiently we can encode the data. Furthermore, there are numerous expressions of general principles that compact or simple representations of data generating processes are “good” ones (Solomonoff, 1964), and numerous researchers have expressed a general view that compression and intelligence are tightly linked (Delétang et al., 2023).
We can also make specific predictions about generalization for different scaling rates of information content capacity. For example, a theorem proving system that proves everything from a fixed set of first principles will exhibit no increase in information content as the set of proven theorems is expanded, and we may surmise that this system will be able to prove many unseen theorems without additional training. A theorem proving system that uses a very large collection of heuristics of limited applicability will exhibit increasing information content requirements as it learns to prove more theorems, and may not be able to prove many unseen theorems.
@@ 1.2.1 Generalization properties of different two-hop algorithms
1.2.2 Empirical evidence regarding two-hop generalization
@@
This evidence is suggestive, but not decisive. The probing results suggest that one aspect of our intuition was sound – that transformers require answers to the first hop to be made available early enough for subsequent processing to be feasible. However, absent proof, there is a large space of possible algorithms that could be used for this purpose. For example, it may be possible in principle to look up the answer to the first hop and encode both this answer and the entire function $f(\cdot,r_{2}):N\to N$ this in activations, which may then be evaluated on $e_{1}$ using an evaluation map which can perhaps be encoded weight-efficiently.
In this work, we assess which of the proposed algorithms best matches the observed information content scaling of transformers trained to answer two-hop questions, which gives us an independent line of evidence regarding which algorithm transformers implement for this task.
@@ 1.3 Summary
  • For one-hop QA, we reproduced previous results of a “knowledge capacity” of about 2 bits per parameter
  • For two-hop QA, we found that the “two function composition” algorithm had a qualitatively better fit to assumptions about knowledge capacity and dataset entropy, but the fit was still questionable
  • We conclude there are significant challenges to applying this method of capacity measurement to understand how transformers learn complex, structured data
@@
    • Two hop QA datasets took about 10 times longer to learn than one-hop QA datasets of similar size
@@ 2.3 Information Content Measurement
$C_{1}\geq E_{1}+|N||A|\mathbb{E}\log_{2}p_{\text{one-hop}}$ (5)
@@
$C_{2}^{\text{independent}}\geq E_{2}^{\text{independent}}-|N||R||A|\mathbb{E}\log_{2}p_{\text{two-hop}}$ (6)
@@
$C_{2}^{\text{recurrent}}\geq$ $E_{2}^{\text{recurrent}}-|N||A|\mathbb{E}\log_{2}p_{\text{eff}}^{\text{recurrent}}$ (7)
$C_{2}^{\text{2f}}\geq$ $E_{2}^{\text{2f}}-|N||A|(\mathbb{E}\log_{2}p_{\text{eff}}^{\text{hop 1}}+\mathbb{E}\log_{2}p_{\text{eff}}^{\text{hop 2}})$ (8)
@@ 3.1 Information Content
Figure 1: Observed information content scaling on one-hop questions for 4 layer transformers. A loss of zero implies that the information content is equal to the dataset entropy, and a loss equal to predicting the uniform distribution over all answers in the dataset yields the information content level represented by “baseline” (this is nontrivial due to needing to learn the set of names in the dataset out of all possible names).
We did not perfectly reproduce the existing result of a 2 bit per parameter information capacity for one-hop questions and answers. Instead, our one-hop training runs yielded an information capacity of around 1.6 bits per parameter (Figure 1). We did successfully replicate the figure in earlier training runs (Figure 8), but these had somewhat different hyperparameters to the two-hop training runs that appear in this paper – they did not employ $\mu$P parametrization and used 4 relations instead of 17 relations.
@@ 3.2 Two Hop Generalization
Figure 4: The generalization gap for two-hop question answering. In most cases the gap increased as the parameter count increased. Models trained on the 1000 profile dataset were an exception, with the largest model on this dataset achieving nearly 0 loss on both the train and evaluation sets. It is possible that even larger models might exhibit smaller generalization gaps on the larger datasets.
@@ 3.4 Discussion and Conclusion
David Johnston came up with the main ideas behind this project, did all experiments and analysis, and wrote nearly all of the paper. Nora Belrose gave advice and guidance throughout the project. David and Nora are funded by a grant from Open Philanthropy. We thank Coreweave for computing resources.
@@ A.1 Two hop recurrent function composition
Next, we consider the case where we have a collection of two-hop questions and answers. We know that these are generated deterministically from a function $f$ mapping one-hop questions to answers. There may be several such functions that lead to the same two-hop questions and answers: if $\forall j:f(e_{i},r_{j})=f(e_{k},r_{j})$ then defining $g$ such that $g(e_{l},r_{m})=f(e_{l},r_{m})$ unless $f(e_{l},r_{m})=e_{i}$, in which case $g(e_{l},r_{m})=e_{k}$ will lead to the same two-hop questions and answers. This is extremely unlikely in our generation scheme - we have $N$ entities, with the probability of a perfect match between any two less than $\frac{1}{N^{1}7}$, which is very small compared to $\frac{1}{C^{N}_{2}}$.
There are unlikely to be other symmetries of $f$ with respect to two-hop question answering. If we want $g(e_{i},r_{j})=e_{k}\neq f(e_{i},r_{j})=e_{p}$ such that $f(e_{p},r_{l})\neq f(e_{k},r_{l})$ for some $l$, then we require $\forall m:g(e_{k},r_{m})=f(e_{k},r_{m})$ (to maintain the same two-hop answers), and also $\forall n:g(e_{q},r_{n})=f(e_{p},r_{n})$ for some $q$ (because $e_{p}$ is an intermediate entity in some two-hop questions). If $e_{q}$ does not have the same relations as $e_{p}$ (which would be very unlikely in our setting), we need an additional entity to play the role of $e_{p}$ and so on. It is very likely that this exploding set of constraints eventually leads to a collision with the requirement to answer two-hop questions correctly: if, for example, $g(e_{k},r_{m})=f(e_{k},r_{m})=e_{q}$ then we require both that $\forall ng(e_{q},r_{n})=f(e_{p},r_{n})$ but also $\forall ng(e_{q},r_{n})=f(e_{q},r_{n})$ to keep the answers to the two-hop questions the same – which, as we’ve already argued, is very unlikely.
@@
Where $e_{11}^{*}=f(e_{1},r_{1})$ and $e_{112}^{*}=f(e_{2},r_{2})$. Let $q_{112}:=P_{2}(e_{112}^{*}|e_{1},r_{1},r_{2})$, $p_{11}:=P_{1}(e_{11}^{*}|e_{1},r_{1})$ and $p_{112}=P_{1}(e_{112}^{*}|e_{11^{*}},r_{2})$. We assume the $p_{ij}$s are mutually independent [^2] The $p_{ij}$s may in fact be negatively correlated because the model is spreading a total budget across all $i$, $j$ pairs, but $|N||R|$ is large and so this correlation will be small..
@@
We make a very crude approximation, valid when $\text{Var}(p)$ is small, that
$\mathbb{E}[\log p_{ij}]$ $\approx\log\left(\frac{-1+\sqrt{1-4|N|(1-|N|e^{\mathbb{E}\log q_{ijk}})}}{2|N|}\right)$

### v2

@@ Abstract
Prior work has found that transformers have an inconsistent ability to learn to answer latent two-hop questions – questions of the form “Who is Bob’s mother’s boss?” We study why this is the case by examining how transformers’ capacity to learn two-hop questions and answers (two-hop QA) scales with their size, motivated by prior work on transformer knowledge capacity for simple fact memorization. We find that capacity scaling and generalization both support the hypothesis that latent two-hop QA requires transformers to learn each fact twice. In contrast, we were unable to find evidence for this conclusing with supervised probing. Capacity scaling for two-hop QA with chain of thought suggests each fact is only learned once, consistent with the better generalization performance of this configuration. We also show that with appropriate dataset parameters, it is possible to “trap” very small models in a regime where they memorize answers to two-hop questions independently, even though they would perform better if they could learn to answer them with function composition. Our findings show that measurement of capacity scaling can complement existing interpretability methods.
@@ 1.1 Motivation
Many existing transformer interpretability methods, like probing (Belinkov, 2022) and sparse autoencoders (Huben et al., 2023) involve extracting interpretable signals from the activations. These allow us to identify features that are relevant to the computation for particular data points, but they also necessarily ignore other parts of the computation. Another ubiquitous method for understanding models is to test how they generalize to situations absent from the training data – this is extremely useful, but so standard it would not typically be considered an “interpretability technique”. Both methods come with limitations: hidden state probing may fail to capture important features of the computation or fail to yield useful hypotheses about how the model is computing answers, and it may be infeasible to test a model’s generalization performance on all tasks of interest.
We investigate a new, complementary method for interpreting transformers: information capacity scaling. Prior work by Allen-Zhu & Li (2024) found that sufficiently trained transformers can memorize facts up to a limit of around 2 bits per parameter, independent of architecture or dataset size.
@@
We hypothesize that the figure of 2 bits per parameter represents how much a transformer trained to saturation can be compressed with respect to a model of computation that matches the transformer’s architecture. This compression will be less efficient than the “ideal” compression in tasks that benefit from recurrence, because recurrent models of computation allow for compact representation of repeated application of functions while transformers must explicitly encode each function application. We will call this imprecise notion of compression the “native information content” of a transformer, and we will study how this figure scales with model size for the task of two-hop question answering.
The native information content of transformers is interesting because, we suggest, it is related to how well transformers generalize. The size of an uncompressed transformer is a very loose upper bound on its information content – an untrained transformer is the same size as a transformer trained to saturation. On the other hand, it is conceivable that a very smart compression algorithm produces a better model of the data generating process than the transformer it is compressing (for example, by exploiting recurrent patterns in the weights of the model being compressed). We speculate that the native information content represents the right level of compression for studying the generalization properties of transformers.
Information compression is closely linked to learning in general. A precise, if narrow, version of this principle is: given a sequence of data $(X_{1},X_{2},...)$ sampled from a distribution $P(X)$, the better we can model $P$, the more efficiently we can encode the data. Many authors have expressed a broader view that compression and intelligvence are tightly linked (Delétang et al., 2023), often partly informed by the classical result of (Solomonoff, 1964) linking representation length with predictive performance.
Practically, we make specific predictions about generalization based on specific algorithmic hypotheses and their information content scaling properties. This rests on the assumption that algorithms with the same information content scaling properties tend to have the same generalization behaviour. While the current work seems to offer modest support for this position, it is not clear how often it is applicable in general. One could also test more general hypotheses; for example, in work that tests whether transformers use correct multihop reasoning or take shortcuts (such as Dziri et al. (2023)) we could measure how information capacity scaling diverges from the scaling of correct multihop reasoning, which may allow us to conclude that transformers employ less efficient algorithms than correct multihop reasoners in general, without offering specific predictions about generalization.
@@ 1.2.1 Generalization properties of different two-hop algorithms
1.2.2 Existing work on multi-hop reasoning in Transformers
@@
In contrast, Yu (2025) using similar probing methods on a dataset of multi-step arithmetic problems found that intermediate steps were often non-recoverable with probing prompted models, even when these models could answer the questions with nontrivial accuracy, though they found fine-tuned models allowed for intermediate step recovery with probing.
Our own probing results are weaker than these: in fact, we failed to recover intermediate answers using probing (see Section D) even when these models had generalization results and capacities that matched the two function composition hypothesis. That is, in our experiments, inferences based on information capacity matched generalization results better than inferences based on probing results.
Transformers are known to be weak at compositional tasks in general. Dziri et al. (2023) found that transformers often failed to learn to generalize to compositional tasks whose graphical representation was not seen in training data. Theoretical analyses have also placed limits on the capabilities of transformers without chain of thought (Merrill et al., 2022; Liu et al., 2023). These limitations are not present for transformers that may use chain of thought reasoning (Pérez et al., 2021).
@@ 1.3 Summary
  • For one-hop QA, we approximately reproduced previous results of a “knowledge capacity” of about 2 bits per parameter, though our measured capacities were somewhat lower
  • For two-hop QA, we found that the “two function composition” algorithm best fit the hypothesis of a 2 bits per parameter knowledge capacity
  • We did not find evidence for two function composition using supervised probing
  • We conclude that knowledge capacity measurement is an alternative method to investigate questions about which algorithms a transformer implements which can be informative where probing fails, but it comes with significant challenges
@@ 2.3 Information Content Measurement
Figure 1: Observed information content scaling on one-hop questions for 4 layer transformers. A loss of zero implies that the information content is equal to the dataset entropy, and a loss equal to predicting the uniform distribution over all answers in the dataset yields the information content level represented by “baseline” (this is nontrivial due to needing to learn the set of names in the dataset out of all possible names). The dashed black line represents a content of 2 bits per parameter.
@@
$C_{1}\geq E_{1}+|N||A|\mathbb{E}\left[\log_{2}p_{\text{one-hop}}\right]$ (5)
@@
$C_{2}^{\text{independent}}\geq E_{2}^{\text{independent}}-|N||R||A|\mathbb{E}\left[\log_{2}p_{\text{two-hop}}\right]$ (6)
@@
$C_{2}^{\text{recurrent}}\geq$ $E_{2}^{\text{recurrent}}-|N||A|\mathbb{E}\left[\log_{2}p_{\text{eff}}^{\text{recurrent}}\right]$ (7)
$C_{2}^{\text{2f}}\geq$ $E_{2}^{\text{2f}}-|N||A|(\mathbb{E}\left[\log_{2}p_{\text{eff}}^{\text{hop 1}}\right]+\mathbb{E}\left[\log_{2}p_{\text{eff}}^{\text{hop 2}}\right])$ (8)
@@ 3.1 Information Content
We did not perfectly reproduce the existing result of a 2 bit per parameter information capacity for one-hop questions and answers. Instead, our one-hop training runs yielded an information capacity of around 1.6 bits per parameter (Figure 1). We did successfully replicate the figure in earlier training runs (Figure 8), but these runs had different hyperparameters to the two-hop training runs that appear in this paper – they did not employ $\mu$P parametrization and used 4 relations instead of 17 relations.
@@ 3.2 Two Hop Generalization
Figure 4: The generalization gap for two-hop question answering. In most cases the gap increased as the parameter count increased. Models trained on the 1000 profile dataset were an exception, with the largest model on this dataset achieving nearly 0 loss on both the train and evaluation sets.
@@ 3.3 Two Hop independent memorization
3.4 Probing
Wang et al. (2024) used a logit lens to show that in their generalising two-hop reasoning circuits, the intermediate entities could be recovered by probing the appropriate token positions. Here we employ supervised linear probes to try to recover intermediate entities at a range of token positions. We expect supervised linear probes to be more effective at extracting linearly retrievable information than the logit lens, which simply applies the model’s unembedding weights to the hidden layer activations without any training on the hidden layer specifically.
Somewhat surprisingly, and in contrast to our capacity investigations, we did not find a signal that strongly differentiated between generalizing and nongeneralizing models. See Tables 2 and 3. In particular, these results show that the intermediate entity is not more retrievable in models where we infer two-function composition than in models where we infer memorization, and in fact the second entity is barely more reterievable than an arbitrary relation of the first entity.
We have compelling evidence of systematic differences between memorizing models and models that implement two function composition – they generalize differently, and exhibit different capacities for learning two hop QA – but these differences were not apparent to supervised probes.
@@ 3.5 Discussion and Conclusion
David Johnston came up with the main ideas behind this project, performed experiments and analysis. Sergio performed additional experiments and contributed analyses. Nora Belrose gave advice and guidance throughout the project. David and Nora are funded by a grant from Open Philanthropy. Sergio contributed as a volunteer. We thank Coreweave for computing resources.
@@ A.1 Two hop recurrent function composition
Consider the case where we have a collection of two-hop questions and answers. We know that these are generated deterministically from a function $f$ mapping one-hop questions to answers. There may be several such functions that lead to the same two-hop questions and answers: if $\forall j:f(e_{i},r_{j})=f(e_{k},r_{j})$ then defining $g$ such that $g(e_{l},r_{m})=f(e_{l},r_{m})$ unless $f(e_{l},r_{m})=e_{i}$, in which case $g(e_{l},r_{m})=e_{k}$ will lead to the same two-hop questions and answers (i.e. if two intermediate entities have identical signatures, then they can be swapped). This is extremely unlikely in our generation scheme - we have $N$ entities, with the probability of a perfect match between any two less than $\frac{1}{N^{17}}$, which is very small compared to $\binom{N}{2}$.
There are unlikely to be other symmetries of $f$ with respect to two-hop question answering. If we want $g(e_{i},r_{j})=e_{k}\neq f(e_{i},r_{j})=e_{p}$ such that $f(e_{p},r_{l})\neq f(e_{k},r_{l})$ for some $l$, then we require $\forall m:g(e_{k},r_{m})=f(e_{k},r_{m})$ (to maintain the same two-hop answers), and also $\forall n:g(e_{q},r_{n})=f(e_{p},r_{n})$ for some $q$ (because $e_{p}$ is an intermediate entity in some two-hop questions). If $e_{q}$ does not have the same relations as $e_{p}$ (which would be very unlikely in our setting), we need an additional entity to play the role of $e_{p}$ and so on. It is very likely that this exploding set of constraints eventually leads to a collision with the requirement to answer two-hop questions correctly: if, for example, $g(e_{k},r_{m})=f(e_{k},r_{m})=e_{q}$ then we require both that $\forall n:g(e_{q},r_{n})=f(e_{p},r_{n})$ but also $\forall n:g(e_{q},r_{n})=f(e_{q},r_{n})$ to keep the answers to the two-hop questions the same – which, as we’ve already argued, is very unlikely.
@@
Where $e_{11}^{*}=f(e_{1},r_{1})$ and $e_{112}^{*}=f(e_{2},r_{2})$. Let $q_{112}:=P_{2}(e_{112}^{*}|e_{1},r_{1},r_{2})$, $p_{11}:=P_{1}(e_{11}^{*}|e_{1},r_{1})$ and $p_{112}=P_{1}(e_{112}^{*}|e_{11^{*}},r_{2})$. We assume the $p_{ij}$s are mutually independent. [^2] The $p_{ij}$s may in fact be negatively correlated because the model is spreading a total budget across all $i$, $j$ pairs, but $|N||R|$ is large and so this correlation will be small.
@@
We make the approximation that
$\mathbb{E}[\log p_{ij}]$ $\approx\log\left(\frac{-1+\sqrt{1-4|N|(1-|N|e^{\mathbb{E}[\log q_{ijk}]})}}{2|N|}\right)$
Empirically, when $\mathbb{E}[p_{ij}]\approx\frac{1}{N}$, $\text{Var}[\log q_{ijk}]\approx 0$ and the approximation is almost exact. On the other hand, if $\mathbb{E}[p_{ij}]]\gg\frac{1}{N}$, then
$\mathbb{E}[q_{ijk}]$ $\approx\mathbb{E}[p_{ij}]^{2}$
$\mathbb{E}[\log p_{ij}]$ $\approx\frac{\mathbb{E}[\log q_{ijk}]}{2}$
$\approx\log\left(\frac{-1+\sqrt{1-4|N|(1-|N|e^{\mathbb{E}\log q_{ijk}})}}{2|N|}\right)$
thus it is a close approximation at both extremes of the values of $p_{ij}$.
@@ Appendix C Additional Information Content Scaling Plots
Appendix D Probing results
Model Details Layer Name Loss Relation Loss Attribute Loss All Loss
Parameters: 890,320 1 5.58 6.67 6.70 5.56
Relations: 4 2 5.56 6.61 6.66 5.26
QA Loss: 3.16 3 5.62 6.53 6.68 5.14
Inferred Algorithm: Memorization 4 5.20 5.75 6.81 4.64
Parameters: 890,320 1 5.57 6.38 6.40 5.57
Relations: 17 2 5.84 6.21 6.40 5.65
QA Loss: 8.18 3 5.94 6.22 6.40 5.76
Inferred Algorithm: 2FC 4 5.68 5.92 6.65 5.46
Parameters: 2,192,400 1 5.26 6.39 6.40 5.27
Relations: 17 2 5.60 6.10 6.40 5.28
QA Loss: 6.56 3 5.69 6.09 6.40 5.34
Inferred Algorithm: 2FC 4 5.20 5.20 6.55 4.57
Parameters: 6,369,424 1 4.57 6.38 6.40 4.57
Relations: 17 2 4.92 5.98 6.23 4.46
QA Loss: 2.25 3 5.02 6.01 6.26 4.59
Inferred Algorithm: 2FC 4 5.05 4.74 6.84 4.28
Parameters: 13,575,328 1 3.98 6.38 6.40 3.99
Relations: 17 2 4.47 5.47 5.98 3.67
QA Loss: 0.19 3 4.69 5.59 6.09 3.92
Inferred Algorithm: 2FC 4 5.89 3.81 6.93 4.32
Table 2: Supervised probe losses for the second entity in two hop questions. An uninformed predictor achieves a loss of 6.47 (this is lower than QA loss as it was only evaluated on the first token of the relevant name). 2FC: two-function composition. Losses refer to probe positions: name loss is the loss of a probe trained on the (initial) name tokens, relation loss for a probe trained on relation tokens, attribute loss for attribute tokens and all loss for a probe trained on all previously mentioned tokens.
Model Details Function Name Loss Relation Loss Attribute Loss All Loss
Parameters: 890,320 1 5.89 6.64 6.67 5.63
Relations: 4 2 5.97 6.62 6.64 5.56
QA Loss: 3.16 3 5.80 6.59 6.63 5.71
Algorithm: Memorization 4 5.36 5.91 6.84 4.89
Parameters: 890,320 1 5.61 6.44 6.46 5.58
Relations: 17 2 5.88 6.26 6.46 5.69
QA Loss: 8.18 3 5.98 6.27 6.45 5.79
Algorithm: 2FC 4 5.73 5.94 6.70 5.49
Parameters: 2,192,400 1 5.31 6.44 6.47 5.30
Relations: 17 2 5.65 6.15 6.40 5.32
QA Loss: 6.56 3 5.74 6.15 6.39 5.37
Algorithm: 2FC 4 5.19 5.26 6.58 4.45
Parameters: 6,369,424 1 4.56 6.43 6.46 4.55
Relations: 17 2 4.94 6.03 6.30 4.46
QA Loss: 2.25 3 5.05 6.06 6.34 4.63
Algorithm: 2FC 4 5.07 4.70 7.05 4.24
Parameters: 13,575,328 1 4.07 6.44 6.46 4.04
Relations: 17 2 4.58 5.60 6.10 3.76
QA Loss: 0.19 3 4.82 5.72 6.21 4.04
Algorithm: 2FC 4 5.81 3.82 7.00 4.29
Table 3: Supervised probe losses for any relation of the first entity in two hop questions. An uninformed predictor achieves a loss of 6.47. 2FC: two-function composition. Losses refer to probe positions: name loss is the loss of a probe trained on the (initial) name tokens, relation loss for a probe trained on relation tokens, attribute loss for attribute tokens and all loss for a probe trained on all previously mentioned tokens.